Solve:
a) logbase5 6-logbase5 x=logbase5 2
b) logbase2 8-logbase2 4=x
c) logbase3 27+ logbase3 9-logbase3 81=logbase3 x
d) logbase2 8+logbase2 4-logbase2 2=x

Answer by stanbon(75887)

Solve:
a) logbase5 6-logbase5 x=logbase5 2
log(6/x) = log2
6/x = 2
x = 3
------------------
b) logbase2 8-logbase2 4=x
log2 (8/4) = x
log2 2 = x
x = 1
-------------------
c) logbase3 27+ logbase3 9-logbase3 81=logbase3 x
3 + 2-4 = x
x = 1
-------------------

d) logbase2 8+logbase2 4-logbase2 2=x
3 + 2 - 1 = x
x = 4
===================
Cheers,
Stan H.